OpenStudy (stephanos100):

HELP!! PLS!! MEDALS! BASIC CALC QUESTION Find an equation of the line tangent to the graph of f at the given point. f(x)=arctan(2x); (1/2, π/4)

OpenStudy (supersmart1001):

the answer is 2x-4y=2-pi

OpenStudy (dumbcow):

First find derivative f'(x) rewrite equation as \[\tan(y) = 2x\] Differentiate \[\sec^2 (y) \frac{dy}{dx} = 2\] \[(1+\tan^2 (y)) \frac{dy}{dx} = 2\] \[\frac{dy}{dx} = \frac{2}{1+ \tan^2 (\tan^{-1} 2x)} = \frac{2}{1+(2x)^2} = \frac{2}{1+4x^2}\] The equation of a line is: \[y = \frac{dy}{dx} (x-\frac{1}{2}) + \frac{\pi}{4}\] plugging in x=1/2 dy/dx = 1 Therefore final answer is: \[y = x-\frac{1}{2} + \frac{\pi}{4}\]